A block slides down on an incline of angle 30° with an acceleration. Find the kinetic friction coeficient

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Kinetic friction coefficient = tan ( theta)

= Tan ( 30)
= ( 1/3)^0.5

Regards
Khimraj
4 years ago
Let the cofficent of kinetic friction be u
( Force Total ) mg / 4 = mg sin Ф - u m g cos Ф 
Solve to get the answer as follows:
sin - u cos = 1 / 4
Substitute : Ф = 30 we get
u = 2 / √3
Please mark the answer as brainliest thank you